OpenStudy (anonymous):

log base 8 (x) + log base 8 (x+4)=k x=?

OpenStudy (cwrw238):

log base 8 x(x+4) = k 8^k = x(x+4)

OpenStudy (cwrw238):

solve this quadratic equation for x